A solution of 49% vinegar is to be mixed with a solution of 32% vinegar to form 153 liters of a 40% solution. How many liters of 49% solution must be used?
Answer by nerdybill(7384)

A solution of 49% vinegar is to be mixed with a solution of 32% vinegar to form 153 liters of a 40% solution. How many liters of 49% solution must be used?
.
Let x = amount (liters) of 49% solution
then
153-x = amount (liters) of 32% solution
.
.49x + .32(153-x) = .40(153)
.49x + 48.96-.32x = 61.2
.17x + 48.96 = 61.2
.17x = 12.24
x = 72 liters